OpenStudy (anonymous):

Which function's domain consists of all real numbers except 3? a) f(x)=3x+3 b) f(x)=3/x-3 c) f(x)=3x^2 d) f(x)=3x-3 e) f(x)= 1/ square root of x-3

OpenStudy (anonymous):

Look at the ones like f(x) = 3/(x-3) is it in brackets? If so, what makes the function undefined/ zero?

OpenStudy (anonymous):

it's f(x)=3/x-3 , it's not in brackets.

OpenStudy (anonymous):

well i mean is it 

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

Yeah, so what happens if you plug in 3?

OpenStudy (anonymous):

so are you saying the answer is B?

OpenStudy (anonymous):

No, I'm asking you a question, what happens when you plug 3 for x?

OpenStudy (anonymous):

The math question is like that, idk how to solve it.

OpenStudy (anonymous):

Domain is the x values, so can you plug every number in this questions and get a solution? It's restricted if you have an undefined or it equals 0.

OpenStudy (anonymous):

So you have f(3) = 3/(3-3) = 3/0 so this can't be all real numbers.

OpenStudy (anonymous):

So what do you suggest is the answer?

OpenStudy (anonymous):

You're looking for the functions that are undefined at x = 3.

OpenStudy (anonymous):

So plug in x in the functions you're given and see if it's undefined, when you divide by 0 you get undefined.

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

You cannot have 0/x or x/0.